### v1

# Globalization of Partial Actions of Ordered Groupoids on Rings

Abstract.
We provide a necessary and sufficient condition to the existence of an ordered globalization of a partial ordered action of an ordered groupoid on a ring and we also present criteria to obtain uniqueness. Furthermore, we apply those results to obtain a Morita context and to show that an inverse semigroup partial action has a globalization (unique up to isomorphism) if, and only if, it is unital.
2020 AMS Subject Classification: Primary. 20L05, 18B40, 20N02. Secondary. 16S35, 20M18
Keywords: ordered groupoid, partial action, ordered globalization, Morita theory, inverse semigroup
1. Introduction
Much has been developed in the theory of actions and partial actions of groupoids. For instance, Galois theories were considered in [4], [5], [8], [14], [24], [25], [26] and [27], actions in sets and duality theorems in [9], generalizations of classic results of group theory such as isomorphism theorems, Lagrange’s Theorem, Sylow theorems, classification of groupoids and properties of cosets in [1], [2], [6] and [21], and partial actions of ordered groupoids in [3], [15] and [22].
The concept of partial group action was given by Exel in [13] aiming to classify a certain class of $C^{*}$-algebras. J. Kellendonk and M. Lawson [19] investigated applications in several other areas, such as graph immersions and inverse semigroups [19], $\mathbb{R}$-trees [28] and tilings of the Euclidean space [18]. Much more can be found in the surveys written by Dokuchaev in [10] and [11]. The definition of partial action of a groupoid was given by Bagio and Paques in [4]. Among other results, they proved a condition for the existence and uniqueness of a globalization. Partial ordered actions of ordered groupoids on rings were introduced in [3], but no study on the existence of globalization has been done so far.
Our main goal in this work is to analyze the existence and uniqueness of ordered globalizations of partial ordered actions of ordered groupoids on rings. In Section 2 we provide a background about partial ordered actions of ordered groupoids on rings and fix some notations. In Section 3 we prove that a partial ordered action of an ordered groupoid on a ring has a globalization if, and only if, it is unital. In Section 4 we define strong partial ordered actions, pseudoassociative groupoids and minimal globalizations, showing that a unital strong partial ordered action of a pseudoassociative groupoid on a ring always has a unique minimal globalization. Finally, in Section 5 we present two applications: we construct a Morita context between the crossed products of an ordered groupoid acting partially on a ring and its globalization and we apply our results in the case of inverse semigroups acting on rings via the Ehresmann-Schein-Nambooripad (ESN) Theorem [20, Theorem 4.1.8], which states that the category of inverse semigroups and premorphisms is isomorphic to the category of inductive groupoids and ordered functors.
Throughout, rings and algebras are associative and unital unless otherwise stated.
2. Preliminaries
We recall that a groupoid $\mathcal{G}$ is a small category in which every morphism is an isomorphism. We denote by $\mathcal{G}_{0}$ the set of objects of $\mathcal{G}$. Observe that $\text{id}:\mathcal{G}_{0}\rightarrow\mathcal{G}$, given by $\text{id}(x)=\text{id}_{x}$, is an injective map and whence we identify $\mathcal{G}_{0}\subseteq\mathcal{G}$. Given $g\in\mathcal{G}$, the domain and the range of $g$ will be denoted by $d(g)$ and $r(g)$, respectively. Hence, $d(g)=g^{-1}g$ and $r(g)=gg^{-1}$. For all $g,h\in\mathcal{G}$, we write $\exists gh$ whenever the product $gh$ is defined. We denote $\mathcal{G}_{2}=\{(g,h)\in\mathcal{G}\times\mathcal{G}:\exists gh\}$.
We say that $\mathcal{G}$ is ordered if there is a partial order $\leq$ on $\mathcal{G}$ such that:
  • (OG1)
    If $g\leq h$ then $g^{-1}\leq h^{-1}$;
  • (OG2)
    For all $g,h,k,\ell\in\mathcal{G}$ such that $g\leq h$, $k\leq\ell$, $\exists gk$ and $\exists h\ell$, we have $gk\leq h\ell$;
  • (OG3)
    Given $g\in\mathcal{G}$ and $e\in\mathcal{G}_{0}$ such that $e\leq d(g)$, there is a unique element $(g|e)\in\mathcal{G}$ which satisfies $(g|e)\leq g$ and $d(g|e)=e$.
  • (OG3*)
    Given $g\in\mathcal{G}$ and $e\in\mathcal{G}_{0}$ such that $e\leq r(g)$, there is a unique element $(e|g)\in\mathcal{G}$ which satisfies $(e|g)\leq g$ and $r(e|g)=e$.
In a partially ordered set $P$, we say that $z\in P$ is the meet of $x\in P$ and $y\in P$ if $z\leq x$, $z\leq y$ and if $w\in P$ is such that $w\leq x,y$ then $w\leq z$. We denote $z=x\wedge y$. A poset where every pair of elements has a meet is called a meet semilattice. If $\mathcal{G}$ is such that $\mathcal{G}_{0}$ is a meet semilattice with respect to $\leq$, we say that $\mathcal{G}$ is an inductive groupoid.
Definition 2.1.
Let $\mathcal{G}$ be an ordered groupoid and $A$ a ring. We say that $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a partial action of $\mathcal{G}$ on $A$ if $A_{r(g)}$ is an ideal of $A$, $A_{g}$ is an ideal of $A_{r(g)}$, $\alpha_{g}:A_{g^{-1}}\to A_{g}$ is an isomorphism of rings, for all $g\in\mathcal{G}$, and
  1. (P1)
    $\alpha_{e}=\text{Id}_{A_{e}}$, for all $e\in\mathcal{G}_{0}$, and $A\subseteq\sum_{e\in\mathcal{G}_{0}}A_{e}$;
  2. (P2)
    $\alpha_{h}^{-1}(A_{g^{-1}}\cap A_{h})\subset A_{(gh)^{-1}}$, for all $(g,h)\in\mathcal{G}_{2}$;
  3. (P3)
    $\alpha_{g}\circ\alpha_{h}(a)=\alpha_{gh}(a)$, for all $(g,h)\in\mathcal{G}_{2},a\in\alpha_{h}^{-1}(A_{g^{-1}}\cap A_{h})$.
The condition $A\subseteq\sum_{e\in\mathcal{G}_{0}}A_{e}$ is required to recover the concept of partial group actions on rings, once if $G$ is a group with identity element $1_{G}$ acting partially on a ring $A$ via $\alpha$, then $A_{1_{G}}=A$ and $\alpha_{1_{G}}=\text{Id}_{A}$. Similar conditions have already appeared in some other papers, such as [9] (on the notion of a “split” $\mathcal{G}$-set) and in (P4) of [23].
By [4, Lemma 1.1], it is also true that:
  1. (i)
    $\alpha_{g}^{-1}=\alpha_{g^{-1}}$, for all $g\in\mathcal{G}$.
  2. (ii)
    $\alpha_{g}(A_{g^{-1}}\cap A_{h})=A_{g}\cap A_{gh}$, for all $(g,h)\in\mathcal{G}_{2}$.
We say that a partial action $\alpha$ is unital if each $A_{g}$ is a unital ring with central and idempotent identity $1_{g}$. Clearly, $A_{g}=A1_{g}=1_{g}A$, for all $g\in\mathcal{G}$. We say that $\alpha$ is preunital if $A_{e}$ is a unital ring with central and idempotent identity $1_{e}$, for all $e\in\mathcal{G}_{0}$.
Furthermore, $\alpha$ is said to be a partial ordered action, or simply P.O. action, if
  1. (PO)
    If $g\leq h$, then $A_{g}\subseteq A_{h}$ and $\alpha_{g}=\alpha_{h}|_{A_{g^{-1}}}$.
Two P.O. actions $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $A$ and $\gamma=(C_{g},\gamma_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $C$ are isomorphic (we denote by $\alpha\simeq\gamma$) if there is a collection of ring isomorphisms $\{\varphi_{e}:A_{e}\to C_{e}:e\in\mathcal{G}_{0}\}$ such that
  1. (i)
    $\varphi_{r(g)}(A_{g})=C_{g}$, for all $g\in\mathcal{G}$,
  2. (ii)
    $\varphi_{r(g)}\circ\alpha_{g}(a)=\gamma_{g}\circ\varphi_{d(g)}(a)$, for all $a\in A_{g^{-1}}$.
A global action of $\mathcal{G}$ on $A$ is a partial action such that $A_{g}=A_{r(g)}$, for all $g\in\mathcal{G}$. Ordered global actions and unital global actions are defined analogously. Of course a global action is unital if and only if it is preunital.
Let $\mathcal{G}$ be an ordered groupoid and $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ a P.O. action of $\mathcal{G}$ on a ring $A$. We define the partial skew groupoid ring $A\ltimes_{\alpha}\mathcal{G}$ by $A\ltimes_{\alpha}\mathcal{G}=\bigoplus_{g\in\mathcal{G}}A_{g}\delta_{g}$, where each $\delta_{g}$ is a symbol. We consider the usual sum and the product is given by
$(a_{g}\delta_{g})(b_{h}\delta_{h})=\begin{cases}\alpha_{g}(\alpha_{g^{-1}}(a_{g})b_{h})\delta_{gh},\text{ if }\exists gh,\\ 0,\text{ otherwise,}\end{cases}$
linearly extended. We have that $A\ltimes_{\alpha}\mathcal{G}$ is a ring but it is not necessarily associative nor unital. A condition to ensure the associativity of $A\ltimes_{\alpha}\mathcal{G}$ is to ask $\alpha$ to be a unital partial action (c.f. [3, Proposition 3.1] and [12, Corollary 3.2]). Also, by [3, Propostion 3.3], if $\mathcal{G}_{0}$ is finite, then $A\ltimes_{\alpha}\mathcal{G}$ is unital with identity $1_{A\ltimes_{\alpha}\mathcal{G}}=\sum_{e\in\mathcal{G}_{0}}1_{e}$.
Definition 2.2.
Let $\mathcal{G}$ be an ordered groupoid and $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ a P.O. action of $\mathcal{G}$ on a ring $A$ such that $A\ltimes_{\alpha}\mathcal{G}$ is associative. Define the ideal $N$ of $A\ltimes_{\alpha}\mathcal{G}$ as $N=\langle a\delta_{g}-a\delta_{h}:g\leq h,a\in A_{g}\subseteq A_{h}\rangle$. We define the partial skew ordered groupoid ring $A\ltimes_{\alpha}^{o}\mathcal{G}$ as $A\ltimes_{\alpha}^{o}\mathcal{G}=\frac{A\ltimes_{\alpha}\mathcal{G}}{N}$.
3. Globalization
From now on, assume that all actions and partial actions are preunital, unless stated otherwise.
Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered action of an ordered groupoid $\mathcal{G}$ on a ring $B$. Let $A$ be a unital ideal of $B$. We shall construct a P.O. action $\alpha$ of $\mathcal{G}$ on $A$ by restricting $\beta$. Consider, for each $e\in\mathcal{G}_{0}$, $A_{e}:=A\cap B_{e}$, which is an ideal of $A$. Now, for all $g\in\mathcal{G}$, define $A_{g}:=A_{r(g)}\cap\beta_{g}(A_{d(g)})$, which is an ideal of $A_{r(g)}$ and also define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, which is a ring isomorphism.
Observe that if $e\leq f$ are elements of $\mathcal{G}_{0}$, then $B_{e}\subseteq B_{f}$, which implies $A_{e}=A\cap B_{e}\subseteq A\cap B_{f}=A_{f}$. Now, if $g\leq h$, then
$A_{g}$ $=A_{r(g)}\cap\beta_{g}(A_{d(g)})\subseteq A_{r(h)}\cap\beta_{g}(A_{d(g)})$
$=A_{r(h)}\cap\beta_{h}(A_{d(g)})\subseteq A_{r(h)}\cap\beta_{h}(A_{d(h)})=A_{h},$
and given arbitrary $a\in A_{g^{-1}}$, $\alpha_{h}(a)=\beta_{h}(a)=\beta_{g}(a)=\alpha_{g}(a)$, so $\alpha_{g}=\alpha_{h}|_{A_{g^{-1}}}$. Hence $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a P.O. action of $\mathcal{G}$ on $A$. We say that $\alpha$ is the standard restriction of $\beta$ to the ideal $A$.
Example 3.1.
Let $\mathcal{G}=\{s,s^{-1},r(s),d(s),e\}$ with $\mathcal{G}_{0}=\{r(s),d(s),e\}$, $e\leq s,$ $and\leq s^{-1}$, $e\leq r(s)$, $e\leq d(s)$.
Let $R$ be a commutative ring and $B=Re_{1}\oplus Re_{2}\oplus Re_{3}$, where $\{e_{1},e_{2},e_{3}\}$ is a set of pairwise orthogonal idempotents whose sum is $1_{B}$. Define
$B_{d(s)}=B_{s^{-1}}=Re_{1}\oplus Re_{2},\quad B_{r(s)}=B_{s}=Re_{2}\oplus Re_{3},\quad B_{e}=Re_{2}$
and $\beta_{g}=Id_{B_{g}}$, for all $g\in\mathcal{G}_{0}$, $\beta_{s}(ae_{1}+be_{2})=be_{2}+ae_{3}$, and $\beta_{s^{-1}}=\beta_{s}^{-1}$.
Let $A=Re_{2}\oplus Re_{3}$. The standard restriction $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ of $\beta$ to $A$ is given by $A_{r(s)}=Re_{2}\oplus Re_{3}$, $A_{d(s)}=Re_{2}=A_{e}$, $A_{s}=Re_{2}=A_{s^{-1}}$, and $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}=Id_{A_{g^{-1}}},\text{ for all }g\in\mathcal{G}$.
Remark 3.2.
Every P.O. action $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $A$ obtained by a standard restriction of an ordered action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$, where $A$ is an ideal of $B$, satisfies the following condition: if $e\leq r(g)$, then $A_{(e|g)}=A_{e}\cap A_{g}$. In fact,
$A_{e}\cap A_{g}$ $=A_{e}\cap A_{r(g)}\cap\beta_{g}(A_{d(g)})=A_{e}\cap\beta_{g}(A_{d(g)})$
$=\beta_{g}(\beta_{g^{-1}}(A_{e}))\cap\beta_{g}(A_{d(g)})=\beta_{g}(A_{d(g)}\cap\beta_{g^{-1}}(A_{e}))$
$=\beta_{g}(A_{d(g)}\cap\beta_{(g^{-1}|e)}(A_{e}))=\beta_{g}(A_{d(g)}\cap B_{r(g^{-1}|e)}\cap\beta_{(g^{-1}|e)}(A_{e}))$
$=\beta_{g}(A_{d(g)}\cap B_{d(e|g)}\cap\beta_{(g^{-1}|e)}(A_{e}))=\beta_{g}(A\cap B_{d(g)}\cap B_{d(e|g)}\cap\beta_{(g^{-1}|e)}(A_{e}))$
$=\beta_{g}(A\cap B_{d(e|g)}\cap\beta_{(g^{-1}|e)}(A_{e}))=\beta_{g}(A_{d(e|g)}\cap\beta_{(g^{-1}|e)}(A_{e}))$
$=\beta_{(e|g)}(A_{d(e|g)}\cap\beta_{(g^{-1}|e)}(A_{e}))=A_{e}\cap\beta_{(e|g)}(A_{d(e|g)})=A_{(e|g)}.$
Now we present a new kind of restriction of $\beta$, which generalizes the standard one. Let $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be an ordered global action of $\mathcal{G}$ on a ring $B$. Consider a unital ideal $A_{e}$ of $B_{e}$, for all $e\in\mathcal{G}_{0}$, such that $A_{e}\subseteq A_{f}$ if $e\leq f$ in $\mathcal{G}_{0}$. Then $A_{g}=A_{r(g)}\cap\beta_{g}(A_{d(g)})$ is an ideal of $A_{r(g)}$, for all $g\in\mathcal{G}$. Define $\alpha_{g}=\beta_{g}|_{A_{g^{-1}}}$, for all $g\in\mathcal{G}$. Consider $A$ the (unital) ideal generated by $\bigcup_{e\in\mathcal{G}_{0}}A_{e}$. Since $A_{e}$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$, and $\beta$ is unital, then $A_{e}$ is an ideal of $B$, for all $e\in\mathcal{G}_{0}$. Hence $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$ and $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a P.O. action of $\mathcal{G}$ on $A$. We say that $\alpha$ is a restriction of $\beta$ to the ring $A$. Every standard restriction of $\beta$ is a restriction of $\beta$.
Consider now the ideal $C_{g}$ of $B$ generated by $\bigcup_{r(h)\leq r(g)}\beta_{h}(A_{d(h)})$. Since $A_{d(h)}$ is unital and $\beta_{h}$ is an isomorphism, for all $h\in\mathcal{G}$, we have that $C_{g}=\sum_{r(h)\leq r(g)}\beta_{h}(A_{d(h)})$. Notice that $C_{g}\subseteq B_{g}$. Defining $\gamma_{g}=\beta_{g}|_{C_{g^{-1}}}$, we have that $\gamma=(C_{g},\gamma_{g})_{g\in\mathcal{G}}$ is an ordered global action of $\mathcal{G}$ on $C=\sum_{e\in\mathcal{G}_{0}}C_{e}$ such that the restriction of $\gamma$ to the ring $A$ coincides with $\alpha$.
The next example illustrates that there are restrictions which are not standard.
Example 3.3.
Let $\mathcal{G}=\{m_{0},m_{1},n_{0},n_{1}\}$, where $\mathcal{G}_{0}=\{m_{0},n_{0}\}$, $m_{1}^{2}=m_{0}$, $n_{1}^{2}=n_{0}$, $m_{0}\leq n_{0}$ and $m_{1}\leq n_{1}$. Let $R$ be a commutative ring and $A=Re_{1}\oplus Re_{2}\oplus Re_{3}\oplus Re_{4}$, where the $e_{i}$’s are central idempotents pairwise orthogonal whose sum is $1_{A}$. Define
$A_{m_{0}}=A_{n_{1}}=A_{n_{0}}=A,\qquad A_{m_{1}}=Re_{1}\oplus Re_{3},$
and
$\alpha_{n_{1}}(ae_{1}+be_{2}+ce_{3}+de_{4})$ $=ce_{1}+de_{2}+ae_{3}+be_{4},$
$\alpha_{m_{1}}(ae_{1}+ce_{3})$ $=ce_{1}+ae_{3},$
[…]
for all $a,b,c,d\in R$.
Then $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a unital P.O. action of $\mathcal{G}$ on $A$. In Theorem 3.5 we will see that $\alpha$ is a restriction of some ordered action $\beta$ of $\mathcal{G}$ on some ring $B$. However, $\alpha$ is not a standard restriction, since $A_{(m_{0}|n_{1})}=A_{m_{1}}\neq A_{n_{1}}=A_{m_{0}}\cap A_{n_{1}}$. Hence there are restrictions that are not standard.
Definition 3.4.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital P.O. action of an ordered groupoid $\mathcal{G}$ on a ring $A$. A unital ordered global action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$ is said to be a globalization of $\alpha$ if for all $e\in\mathcal{G}_{0}$ there is an injective homomorphism of rings $\varphi_{e}:A_{e}\to B_{e}$ such that
  1. (i)
    $\varphi_{e}(A_{e})$ is an ideal of $B_{e}$;
  2. (ii)
    $\varphi_{r(g)}(A_{g})=\varphi_{r(g)}(A_{r(g)})\cap\beta_{g}(\varphi_{d(g)}(A_{d(g)}))$, for all $g\in\mathcal{G}$;
  3. (iii)
    $\beta_{g}\circ\varphi_{d(g)}(a)=\varphi_{r(g)}\circ\alpha_{g}(a)$, for all $a\in A_{g^{-1}}$ and $g\in\mathcal{G}$;
  4. (iv)
    $B_{g}=\sum_{r(h)\leq r(g)}\beta_{h}\left(\varphi_{d(h)}(A_{d(h)})\right)$.
Now we shall prove that a P.O. action has a globalization if, and only if, it is unital.
Theorem 3.5.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital P.O. action of an ordered groupoid $\mathcal{G}$ on a ring $A$. Then $\alpha$ has a globalization $\beta$ if and only if $\alpha$ is unital.
Proof.
($\Rightarrow$): Once $\varphi_{r(g)}(A_{g})=\varphi_{r(g)}(A_{r(g)})\cap\beta_{g}(\varphi_{d(g)}(A_{d(g)}))$ and $A_{r(g)}$ and $A_{d(g)}$ are unital, the assertion clearly holds.
($\Leftarrow$): Assume that every $A_{g}$ is unital with identity $1_{g}$. Let $\mathcal{F}:=\mathcal{F}(\mathcal{G},A)$ be the ring of all maps $\mathcal{G}\to A$ and define $\mathcal{G}_{g}=\{h\in\mathcal{G}:r(h)\leq r(g)\}$.
Take $F_{g}=\{f\in\mathcal{F}:f(h)=0,\text{ for all }h\notin\mathcal{G}_{g}\}$. Since $\mathcal{G}_{g}=\mathcal{G}_{r(g)}$, it follows that $F_{g}=F_{r(g)}$. If $h\leq g$, then $r(h)\leq r(g)$, so $F_{h}\subseteq F_{g}$. Clearly $F_{g}$ is an ideal of $\mathcal{F}$, for all $g\in\mathcal{G}$.
Now fix the notation $f(h):=f|_{h}$, for all $f\in\mathcal{F}$ and $h\in\mathcal{G}$.
For $g\in\mathcal{G}$ and $f\in F_{g^{-1}}$, define $\gamma_{g}:F_{g^{-1}}\to F_{g}$ by
$\gamma_{g}(f)|_{h}=\begin{cases}f((g^{-1}|r(h))h),\text{ if }h\in\mathcal{G}_{g},\\ 0,\text{ otherwise.}\end{cases}$
Notice that if $h\in\mathcal{G}_{g}$, then $r(h)\leq r(g)=d(g^{-1})$, and hence the restriction $(g^{-1}|r(h))$ is defined, as well as the product $(g^{-1}|r(h))h$. Moreover, $r((g^{-1}|r(h))h)=r(g^{-1}|r(h))\leq r(g^{-1})$, from where it follows that $\gamma_{g}$ is well defined. It is easy to see that $\gamma_{g}$ is a ring homomorphism.
For $g\in\mathcal{G}$, $f\in F_{g^{-1}}$ and $h\in\mathcal{G}_{g^{-1}}$, we have that
$\gamma_{g^{-1}}\circ\gamma_{g}(f)|_{h}$ $=\gamma_{g}(f)|_{(g|r(h))h}=f((g^{-1}|r(g|r(h)))(g|r(h))h)$
=(*$\stackrel{{\scriptstyle(*)}}{{=}}f((g^{-1}g|r(h))h)=f((d(g)|r(h))h)\stackrel{{\scriptstyle(**)}}{{=}}f(r(h)h)=f(h),$
where the equality (***) follows by [20, Proposition 4.1.3] and the equality (***** *) follows since $r(h)\leq d(g)$. Similarly $\gamma_{g}\circ\gamma_{g^{-1}}(f)|_{h}=f(h)$. Thus we have that $\gamma_{g}$ is a ring isomorphism.
Observe that $\gamma_{e}=I_{F_{e}}$, for all $e\in\mathcal{G}_{0}$ and
$\gamma_{gh}(f)|_{k}$ $=f((h^{-1}g^{-1}|r(k))k)\stackrel{{\scriptstyle(*)}}{{=}}f((h^{-1}|r(g^{-1}|r(k)))(g^{-1}|r(k))k)=\gamma_{g}\circ\gamma_{h}(f)|_{k},$
for all $f\in F_{h^{-1}}$, $k\in\mathcal{G}_{g}$ and $(g,h)\in\mathcal{G}_{2}$, where the equality (***) follows again by [20, Proposition 4.1.3]. Then $\gamma=(F_{g},\gamma_{g})_{g\in\mathcal{G}}$ is a global action of $\mathcal{G}$ on $\mathcal{F}$.
To see that $\gamma$ is ordered, observe that if $h\leq g$ in $\mathcal{G}$ and $f\in F_{h^{-1}}$, then
$\gamma_{g}(f)|_{k}$ $=f((g^{-1}|r(k))k)\stackrel{{\scriptstyle(*)}}{{=}}f((h^{-1}|r(k))k)=\gamma_{h}(f)|_{k},$
that is, $\gamma_{h}=\gamma_{g}|_{F_{h^{-1}}}$. The equality (*)(*)( * ) follows from the uniqueness of restrictions, because $(h^{-1}|r(k))\leq h^{-1}\leq g^{-1}$ and $d(h^{-1}|r(k))=r(k)$. Hence $\gamma$ is an ordered global action of $\mathcal{G}$ on $\mathcal{F}$.
Now, for all $e\in\mathcal{G}_{0}$, define $\varphi_{e}:A_{e}\to F_{e}$ by
$\varphi_{e}(a)|_{h}=\begin{cases}\alpha_{h^{-1}}(a1_{h}),\text{ if }r(h)=e,\\ 0,\text{ otherwise,}\end{cases}$
for all $a\in A_{e},h\in\mathcal{G}$. We have that $\varphi_{e}(a)|_{e}=\alpha_{e}(a1_{e})=a$, thus $\varphi_{e}$ is a monomorphism, for all $e\in\mathcal{G}_{0}$.
Let $B_{g}$ be the ideal generated by $\bigcup_{r(h)\leq r(g)}\gamma_{h}(\varphi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}=\left\langle\bigcup_{e\in\mathcal{G}_{0}}B_{e}\right\rangle$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$. By construction, $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ is an ordered global action of $\mathcal{G}$ on $B$. The next step is to show that $\beta$ is a globalization of $\alpha$.
We will start by verifying (iii). Let $g\in\mathcal{G}$, $a\in A_{g^{-1}}$ and $h\in\mathcal{G}$. If $r(h)=r(g)$, then
$\varphi_{r(g)}(\alpha_{g}(a))|_{h}=\alpha_{h^{-1}}(\alpha_{g}(a)1_{h})$
and
$\beta_{g}(\varphi_{d(g)}(a))|_{h}=\varphi_{d(g)}(a)|_{g^{-1}h}=\alpha_{h^{-1}g}(a1_{g^{-1}h}).$
Since $a\in A_{g^{-1}}$, $a1_{g^{-1}h}\in A_{g^{-1}}\cap A_{g^{-1}h}$. Thus
$\beta_{g}(\varphi_{d(g)}(a))$ $=\alpha_{h^{-1}g}(a1_{g^{-1}h})=\alpha_{h^{-1}}(\alpha_{g}(a1_{g^{-1}h}1_{g^{-1}}))$
$=\alpha_{h^{-1}}(\alpha_{g}(a1_{g^{-1}})\alpha_{g}(1_{g^{-1}h}1_{g^{-1}}))=\alpha_{h^{-1}}(\alpha_{g}(a1_{g^{-1}})1_{g}1_{h})$
$=\alpha_{h^{-1}}(\alpha_{g}(a1_{g^{-1}})1_{h}),$
because $\alpha_{g}(A_{g^{-1}}\cap A_{g^{-1}h})=A_{g}\cap A_{h}$ implies $\alpha_{g}(1_{g^{-1}h}1_{g^{-1}})=1_{g}1_{h}$.
On the one hand, if $r(g)\neq r(h)$, we have that $\varphi_{r(g)}(\alpha_{g}(a))|_{h}=0$. On the other hand, if $r(h)<r(g)$,
$\beta_{g}(\varphi_{d(g)}(a))|_{h}=\varphi_{d(g)}(a)|_{(g^{-1}|r(h))h}=\begin{cases}\alpha_{h^{-1}(r(h)|g)}(a1_{(g^{-1}|r(h))h}),\text{ if }r(g^{-1}|r(h))=d(g),\\ 0,\text{ otherwise.}\end{cases}$
Notice that if $r(g^{-1}|r(h))=d(g)$, then since $g^{-1}\leq g^{-1}$ and $r(g^{-1})=d(g)$, from the uniqueness of corestrictions, we have that $g^{-1}=(d(g)|g^{-1})=(d(g)|(g^{-1}|r(h)))=(g^{-1}|r(h))$. But this implies that $r(g)=d(g^{-1})=d(g^{-1}|r(h))=r(h)$, which is a contradiction. Therefore $d(g^{-1}|r(h))<d(g^{-1})$. Thus $\beta_{g}(\varphi_{d(g)}(a))|_{h}=0$. If $r(g)<r(h)$ or $r(g)$ and $r(h)$ are incomparable, this equality also holds, showing (iii).
Now we proceed to (ii). Let $g\in\mathcal{G}$ and $c\in\varphi_{r(g)}(A_{r(g)})\cap\beta_{g}(\varphi_{d(g)}(A_{d(g)}))$. Then there are $a\in A_{r(g)}$, $b\in A_{d(g)}$ such that $c=\varphi_{r(g)}(a)=\beta_{g}(\varphi_{d(g)}(b))$. Therefore,
$a=\alpha_{r(g)^{-1}}(a1_{r(g)})=\varphi_{r(g)}(a)|_{r(g)}=\beta_{g}(\varphi_{d(g)}(b))|_{r(g)}=\varphi_{d(g)}(b)|_{g^{-1}}=\alpha_{g}(b1_{g^{-1}})\in A_{g},$
from where $c=\varphi_{r(g)}(a)\in\varphi_{r(g)}(A_{g})$.
Conversely, if $c\in\varphi_{r(g)}(A_{g})$, then $c=\varphi_{r(g)}(a)$, for some $a\in A_{g}$. Taking $b=\alpha_{g^{-1}}(a)\in A_{g^{-1}}$, we have by (iii) that $\beta_{g}(\varphi_{d(g)}(b))=\varphi_{r(g)}(\alpha_{g}(b))=\varphi_{r(g)}(a)$. Hence $c\in\varphi_{r(g)}(A_{g})\cap\beta_{g}(\varphi_{d(g)}(A_{g^{-1}}))\subseteq\varphi_{r(g)}(A_{r(g)})\cap\beta_{g}(\varphi_{d(g)}(A_{d(g)}))$.
It only remains for us to show that $\varphi_{e}(A_{e})$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$. For this, it is sufficient to show that $\varphi_{e}(b)\beta_{h}(\varphi_{d(h)}(a)),\beta_{h}(\varphi_{d(h)}(a))\varphi_{e}(b)\in\varphi_{e}(A_{e})$, for all $h\in\mathcal{G}_{e},a\in A_{d(h)}$ and $b\in A_{e}$. Given $k\in\mathcal{G}$, we have two possibilities:
Case (1): $r(k)\leq r(h)\leq e$.
Observe that
$\beta_{h}(\varphi_{d(h)}(a))\varphi_{e}(b)|_{k}=\beta_{h}(\varphi_{d(h)}(a))|_{k}\varphi_{e}(b)|_{k}=\varphi_{d(h)}(a)|_{(h^{-1}|r(k))k}\varphi_{e}(b)|_{k}.$
We have that
$\varphi_{e}(b)|_{k}=\begin{cases}\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=e,\\ 0,\text{ otherwise.}\end{cases}$
Since $r(k)=e$ and $r(k)\leq r(h)\leq e$ imply $r(k)=r(h)=e$, we obtain
$\beta_{h}(\varphi_{d(h)}(a))\varphi_{e}(b)|_{k}$ =$=\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}k})\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}k})1_{k^{-1}h}1_{k^{-1}}\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
=(*)$\stackrel{{\scriptstyle(*)}}{{=}}\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}k})\alpha_{k^{-1}h}(1_{h^{-1}k}1_{h^{-1}})\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\alpha_{k^{-1}h}(a1_{h^{-1}}1_{h^{-1}k})\alpha_{k^{-1}}(b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\alpha_{k^{-1}}(\alpha_{h}(a1_{h^{-1}})b1_{k}),\text{ if }r(k)=r(h)=e,\\ 0,\text{ otherwise}\end{cases}$
$=\varphi_{e}(\alpha_{h}(a1_{h^{-1}})b)|_{k},$
where (*)(*)( * ) holds because Proposition (i)(ii) implies that $\alpha_{k^{-1}h}(A_{h^{-1}k}\cap A_{h^{-1}})=A_{k^{-1}h}\cap A_{k^{-1}}$.
Case (2): $r(h)\leq r(k)\leq e$ or $r(k)$ and $r(h)$ incomparable.
In this case $\beta_{h}(\varphi_{d(h)}(a))\varphi_{e}(b)|_{k}=\beta_{h}(\varphi_{d(h)}(a))|_{k}\varphi_{e}(b)|_{k}=0\cdot\varphi_{e}(b)|_{k}=0$ directly from the definition of $\beta$.
The verification that $\varphi_{e}(b)\beta_{h}(\varphi_{d(h)}(a))\in\varphi_{e}(A_{e})$ is similar, and then $\varphi_{e}(A_{e})$ is an ideal of $B_{e}$. ∎
Example 3.6.
Let $\mathcal{G}$, $R$, $B$ and $A$ be as in Example 3.1. We will construct a globalization for $\alpha$. Consider $\mathcal{F}=\mathcal{F}(\mathcal{G},A)$. We have that $\mathcal{F}\simeq A^{5}$ via $f\in\mathcal{F}\mapsto(f(s),f(s^{-1}),f(r(s)),f(d(s)),$ $f(e))$.
Since $\mathcal{G}_{s}=\mathcal{G}_{r(s)}=\{s,r(s),e\}$, $\mathcal{G}_{s^{-1}}=\mathcal{G}_{d(s)}=\{s^{-1},d(s),e\}$ and $\mathcal{G}_{e}=\{e\}$, we have that
$F_{s}$ =$=\{f\in\mathcal{F}:f(g)=0,\text{ for all }g\notin\mathcal{G}_{s}\}\simeq A\times 0\times A\times 0\times A,$
$F_{s^{-1}}$ =$=\{f\in\mathcal{F}:f(g)=0,\text{ for all }g\notin\mathcal{G}_{s^{-1}}\}\simeq 0\times A\times 0\times A\times A,$
$F_{e}$ =$=\{f\in\mathcal{F}:f(g)=0,\text{ for all }g\notin\mathcal{G}_{e}\}\simeq 0\times 0\times 0\times 0\times A.$
Thus, $\gamma_{s}(0,a,0,b,c)=(b,0,a,0,c)$, $\gamma_{s^{-1}}(a,0,b,0,c)=(0,b,0,a,c)$ and $\gamma_{g}=Id_{F_{g}}$, for all $g\in\mathcal{G}_{0}$. Furthermore, notice that $\varphi_{r(s)}(ae_{2}+be_{3})=(ae_{2},0,ae_{2}+be_{3},0,0)$, $\varphi_{d(s)}(ae_{2})=(0,ae_{2},0,ae_{2},0)$ and $\varphi_{e}(ae_{2})=(0,0,0,0,ae_{2})$.
Given $ae_{2}+be_{3}\in A_{r(s)}$, $ce_{2}\in A_{d(s)}$, $de_{2}\in A_{e}$, we obtain
$\gamma_{r(s)}(\varphi_{r(s)}(ae_{2}+be_{3}))+\gamma_{s}(\varphi_{d(s)}(ce_{2}))+\gamma_{e}(\varphi_{e}(de_{2}))$
=\;= $\gamma_{r(s)}(ae_{2},0,ae_{2}+be_{3},0,0)+\gamma_{s}(0,ce_{2},0,ce_{2},0)+\gamma_{e}(0,0,0,0,de_{2})$
=\;= $(ae_{2},0,ae_{2}+be_{3},0,0)+(ce_{2},0,ce_{2},0,0)+(0,0,0,0,de_{2})$
=\;= $((a+c)e_{2},0,(a+c)e_{2}+be_{3},0,de_{2}),$
from where it follows that $B_{s}^{\prime}=\{(ae_{2},0,ae_{2}+be_{3},0,ce_{2}):a,b,c\in R\}$.
Similarly, $B_{s^{-1}}^{\prime}=\{(0,ae_{2}+be_{3},0,ae_{2},0,ce_{2}):a,b,c\in R\}$ and $B_{e}^{\prime}=Re_{2}(0,0,0,0,1)$, thus $B^{\prime}=B_{s}^{\prime}+B_{s^{-1}}^{\prime}+B_{e}^{\prime}=\{(ae_{2},be_{2}+ce_{3},ae_{2}+de_{3},be_{2},fe_{3}):a,b,c,d,f\in R\}$.
Then $\beta^{\prime}=(B_{g}^{\prime},\beta_{g}^{\prime})_{g\in\mathcal{G}}$ is a globalization of $\alpha$. Observe that $\beta$ is also a globalization of $\alpha$, taking $\varphi_{g}=\iota_{g}:A_{g}\to B_{g}$ as the usual inclusion $A_{g}\mapsto A\cap B_{g}$, for all $g\in\mathcal{G}_{0}$. However $\beta^{\prime}\not\simeq\beta$, since $B_{s}\simeq R^{2}\not\simeq R^{3}\simeq B_{s}^{\prime}$.
4. Uniqueness
In this section we will give a sufficient condition for a P.O. action to have a unique globalization. In order to achieve this goal, we introduce the notion of a minimal globalization.
Definition 4.1.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital P.O. action of $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. We say that $\beta$ is a minimal globalization if
  1. (iv’)
    $B_{g}=\sum_{r(h)=r(g)}\beta_{h}(\varphi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$,
holds.
Moreover, we introduce a specific type of P.O. action. Given an ordered groupoid $\mathcal{G}$, we define the pseudoproduct *** as
$g*h=\begin{cases}(g|d(g)\wedge r(h))(d(g)\wedge r(h)|h),\text{ if }d(g)\wedge r(h)\text{ is defined},\\ \text{undefined, otherwise.}\end{cases}$
Definition 4.2.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a P.O. action of $\mathcal{G}$ on a ring $A$. We say that $\alpha$ is a strong P.O. action [16, Definition 3.7] if
  1. (PS)
    $\alpha_{g}\circ\alpha_{h}=\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}}$, for all $g,h\in\mathcal{G}$ such that $\exists g*h$.
Remark 4.3.
Since (PS) holds, notice that $\text{dom}(\alpha_{g}\circ\alpha_{h})=\text{dom}(\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}})$, for all $g,h\in\mathcal{G}$ such that $\exists g*h$. But this is the same as saying that $\alpha_{g}(A_{g^{-1}}\cap A_{h})=A_{g}\cap A_{g*h}$.
In the case where the action is global, we have that strong global ordered actions are precisely global ordered actions. In the case where $\mathcal{G}$ is trivially ordered with the equality, the concepts of strong P.O. action and partial action coincide.
Lemma 4.4.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a strong P.O. action of $\mathcal{G}$ on a ring $A$. The following statements are true:
  1. (i)
    If $e\leq r(g)$, $A_{(e|g)}=A_{e}\cap A_{g}$.
  2. (ii)
    If $e\leq d(g)$, then $A_{(g|e)}=A_{g}\cap A_{r(g|e)}$.
  3. (iii)
    If $e,f\in\mathcal{G}_{0}$ are such that $e\wedge f$ is defined, then $A_{e\wedge f}=A_{e}\cap A_{f}$.
Proof.
(i): If $e\leq r(g)$, then $\alpha_{e}\circ\alpha_{g}=\alpha_{e*g}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}$.
Now $A_{(e|g)}=\text{Im}(\alpha_{(e|g)})=\text{Im}(\alpha_{e}\circ\alpha_{g})=\alpha_{e}(A_{e}\cap A_{g})=A_{e}\cap A_{g}$.
(ii): Notice that $(g|e)=(r(g|e)|g)$ by the uniqueness of restrictions. Hence $A_{(g|e)}=A_{(r(g|e)|g)}=A_{g}\cap A_{r(g|e)}$.
(iii): From (PS) it follows that $\alpha_{e}\circ\alpha_{f}=\alpha_{e*f}\circ\alpha_{f}=\alpha_{e\wedge f}\circ\alpha_{f}=\alpha_{e\wedge f}$. Therefore $A_{e\wedge f}=\text{Im}(\alpha_{e\wedge f})=\text{Im}(\alpha_{e}\circ\alpha_{f})=\alpha_{e}(A_{e}\cap A_{f})=A_{e}\cap A_{f}.$
Definition 4.5.
Let $\mathcal{G}$ be an ordered groupoid. We say that $\mathcal{G}$ is pseudoassociative if $\exists(g*h)*k\iff\exists g*(h*k)$, for all $g,h,k\in\mathcal{G}$.
Remark 4.6.
In the case that $\exists(g*h)*k$ and $\exists g*(h*k)$, we have by [20, Lemma 4.1.6] that these pseudoproducts coincide. Inductive groupoids are always pseudoassociative.
Theorem 4.7.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a strong P.O. action of a pseudoassociative groupoid $\mathcal{G}$ on a ring $A$. Then $\alpha$ admits a minimal globalization if, and only if, $\alpha$ is unital. Furthermore, the minimal globalization is unique up to isomorphisms.
Proof.
Consider, as in the proof of Theorem 3.5, $\mathcal{F}=\mathcal{F}(\mathcal{G},A)$ the ring of the maps of $\mathcal{G}$ on $A$. Given $g\in\mathcal{G}$, define $\mathcal{E}_{g}=\{h\in\mathcal{G}:\exists g^{-1}*h\}$. Define $F_{g}=\{f\in\mathcal{F}:f(h)=0,\text{ for all }h\notin\mathcal{E}_{g}\}$.
Clearly $F_{g}=F_{r(g)}$, for all $g\in\mathcal{G}$. Consider $\gamma_{g}:F_{g^{-1}}\to F_{g}$, where
$\gamma_{g}(f)|_{h}=\begin{cases}f(g^{-1}*h),\text{ if }h\in\mathcal{E}_{g},\\ 0,\text{ otherwise.}\end{cases}$
It is easy to see that $\gamma_{g}$ is a well defined ring homomorphism, for all $g\in\mathcal{G}$. If $(g,h)\in\mathcal{G}_{2}$, $f\in F_{h^{-1}}$ and $k\in\mathcal{E}_{g}$ then
$\gamma_{gh}(f)|_{k}$ $=f((gh)^{-1}*k)=f((h^{-1}g^{-1})*k)$
$=f((h^{-1}*g^{-1})*k)=f(h^{-1}*(g^{-1}*k))$
$=\gamma_{h}(f)|_{g^{-1}*k}=\gamma_{g}\circ\gamma_{h}(f)|_{k}.$
If $k\notin\mathcal{E}_{g}$, then $\gamma_{gh}(f)|_{k}=0=\gamma_{g}\circ\gamma_{h}(f)|_{k}$, so $\gamma_{gh}=\gamma_{g}\circ\gamma_{h}$.
For all $e\in\mathcal{G}_{0}$, define $\psi_{e}:A_{e}\to F_{e}$ by
$\psi_{e}(a)|_{h}=\begin{cases}\alpha_{h^{-1}}(a1_{h}),\text{ if }h\in\mathcal{E}_{e},\\ 0,\text{ otherwise},\end{cases}$
for all $a\in A_{e},h\in\mathcal{G}$. We have that $\psi_{e}(a)|_{e}=\alpha_{e}(a1_{e})=a$, for all $a\in A_{e}$. Hence $\psi_{e}$ is a monomorphism, for all $e\in\mathcal{G}_{0}$.
Let $B_{g}$ be the ideal generated by $\bigcup_{r(h)=r(g)}\gamma_{h}(\psi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$.
Given $e\in\mathcal{G}_{0}$, $g\in\mathcal{G}$ with $r(g)=e$ and $a\in A_{d(g)}$, we have that
$\beta_{e}\circ\beta_{g}(\psi_{d(g)}(a))=\gamma_{e}\circ\gamma_{g}(\psi_{d(g)}(a))=\gamma_{eg}(\psi_{d(g)}(a))=\gamma_{g}(\psi_{d(g)}(a))=\beta_{g}(\psi_{d(g)}(a)),$
from where it follows that $\beta_{e}=\text{Id}_{B_{e}}$. Since $\gamma_{g}\circ\gamma_{h}=\gamma_{gh}$, for all $(g,h)\in\mathcal{G}_{2}$, then $\beta_{g}\circ\beta_{h}=\beta_{gh}$, for all $(g,h)\in\mathcal{G}_{2}$.
Now we prove that $\beta$ satisfies (PS). Indeed, given $g,h,k,\ell\in\mathcal{G}$ such that $h\in\mathcal{E}_{g}$, $r(k)=d(h)$ and $a\in A_{d(k)}$, we have that
$\beta_{g}\circ\beta_{h}(\beta_{k}(\psi_{d(k)}(a)))|_{\ell}$ =$=\begin{cases}\beta_{h}(\beta_{k}(\psi_{d(k)}(a)))|_{g^{-1}*\ell},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\beta_{k}(\psi_{d(k)}(a))|_{h^{-1}*(g^{-1}*\ell)},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\beta_{k}(\psi_{d(k)}(a))|_{(h^{-1}*g^{-1})*\ell},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\beta_{k}(\psi_{d(k)}(a))|_{((d(h)*h^{-1})*g^{-1})*\ell},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\beta_{k}(\psi_{d(k)}(a))|_{d(h)*(h^{-1}*g^{-1})*\ell},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
=$=\begin{cases}\beta_{d(h)}(\beta_{k}(\psi_{d(k)}(a)))|_{(h^{-1}*g^{-1})*\ell},\text{ if }\ell\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
$=\beta_{g*h}\circ\beta_{d(h)}(\beta_{k}(\psi_{d(k)}(a)))|_{\ell}.$
To finish the proof, we shall show that $\beta$ satisfies (i)-(iii) from the definition of minimal globalization.
(iii): Let $g,h\in\mathcal{G}$ e $a\in A_{g^{-1}}$. Then
$\psi_{r(g)}\circ\alpha_{g}(a)|_{h}$ =$=\begin{cases}\alpha_{h^{-1}}(\alpha_{g}(a)1_{h}),\text{ if }h\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}$
and
$\beta_{g}\circ\psi_{d(g)}(a)|_{h}$ =$=\begin{cases}\psi_{d(g)}(a)|_{g^{-1}*h},\text{ if }h\in\mathcal{E}_{g},\\ 0,\text{ otherwise}\end{cases}=\begin{cases}\alpha_{h^{-1}*g}(a1_{g^{-1}*h}),\text{ if }h\in\mathcal{E}_{g},\\ 0,\text{ otherwise.}\end{cases}$
Then we can work only with the case in which the meet $e:=r(g)\wedge r(h)$ is defined. As $a\in A_{g^{-1}}$ we have that $a1_{g^{-1}*h}\in A_{g^{-1}}\cap A_{g^{-1}*h}=A_{(e|g^{-1})}\cap A_{g^{-1}*h}$. Hence
$\beta_{g}(\psi_{d(g)}(a))|_{h}$ $=\alpha_{h^{-1}*g}(a1_{g^{-1}*h})=\alpha_{(h^{-1}|e)}\circ\alpha_{(e|g)}(a1_{g^{-1}*h})$
$=\alpha_{(h^{-1}|e)}(\alpha_{(e|g)}(a1_{(g^{-1}|e)})1_{(e|h)})=\alpha_{h^{-1}}(\alpha_{g}(a1_{(g^{-1}|e)})1_{(e|h)})$
$=\alpha_{h^{-1}}(\alpha_{g}(a)\alpha_{g}(1_{g^{-1}}1_{(g^{-1}|e)})1_{(e|h)})=\alpha_{h^{-1}}(\alpha_{g}(a)\alpha_{(e|g)}(1_{(g^{-1}|e)})1_{(e|h)})$
$=\alpha_{h^{-1}}(\alpha_{g}(a)1_{(e|g)}1_{(e|h)})=\alpha_{h^{-1}}(\alpha_{g}(a)1_{e}1_{g}1_{h})$
$=\alpha_{h^{-1}}(\alpha_{g}(a)1_{r(g)}1_{r(h)}1_{g}1_{h})=\alpha_{h^{-1}}(\alpha_{g}(a)1_{g}1_{h})=\alpha_{h^{-1}}(\alpha_{g}(a)1_{h}).$
Therefore $\beta_{g}\circ\psi_{d(g)}(a)=\psi_{r(g)}\circ\alpha_{g}(a)$.
(ii): Let $g\in\mathcal{G}$ and $c\in\psi_{r(g)}(A_{r(g)})\cap\beta_{g}(\psi_{d(g)}(A_{d(g)}))$. Then there are $a\in A_{r(g)},b\in A_{d(g)}$ such that $c=\psi_{r(g)}(a)=\beta_{g}(\psi_{d(g)}(b))$. Thus,
$a=\alpha_{r(g)^{-1}}(a1_{r(g)})=\psi_{r(g)}(a)|_{r(g)}=\beta_{g}(\psi_{d(g)}(b))|_{r(g)}=\psi_{d(g)}(b)|_{g^{-1}}=\alpha_{g}(b1_{g^{-1}})\in A_{g},$
from where it follows that $c=\psi_{r(g)}(a)\in\psi_{r(g)}(A_{g})$.
Conversely, if $c\in\psi_{r(g)}(A_{g})$, then $c=\psi_{r(g)}(a)$, for some $a\in A_{g}$. Taking $b=\alpha_{g^{-1}}(a)\in A_{g^{-1}}$, we have from (iii) that $\beta_{g}(\psi_{d(g)}(b))=\psi_{r(g)}(\alpha_{g}(b))=\psi_{r(g)}(a)=c$.
Thus $c\in\psi_{r(g)}(A_{g})\cap\beta_{g}(\psi_{d(g)}(A_{g^{-1}}))\subseteq\psi_{r(g)}(A_{r(g)})\cap\beta_{g}(\psi_{d(g)}(A_{d(g)}))$.
(i): Let $e\in\mathcal{G}_{0}$, $h\in\mathcal{G}$ be such that $r(h)=e$, $a\in A_{d(h)}$ and $b\in A_{e}$. It is enough to prove that $\psi_{e}(b)\beta_{h}(\psi_{d(h)}(a)),\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)\in\psi_{e}(A_{e})$. For that, consider $k\in\mathcal{G}$. We have that $\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}=\beta_{h}(\psi_{d(h)}(a))|_{k}\psi_{e}(b)|_{k}$. On the one hand,
$\psi_{e}(b)|_{k}=\begin{cases}\alpha_{k^{-1}}(b1_{k}),\text{ if }k\in\mathcal{E}_{e},\\ 0,\text{ otherwise.}\end{cases}$
On the other hand,
$\beta_{h}(\psi_{d(h)}(a))|_{k}$ =$=\begin{cases}\psi_{d(h)}(a)|_{h^{-1}*k},\text{ if }k\in\mathcal{E}_{h}=\mathcal{E}_{e},\\ 0,\text{ otherwise}\end{cases}=\begin{cases}\alpha_{k^{-1}*h}(a1_{h^{-1}*k}),\text{ if }k\in\mathcal{E}_{e},\\ 0,\text{ otherwise.}\end{cases}$
Thus we can work only with the case in which $k\in\mathcal{E}_{e}$. Then, denoting by $f=r(k)\wedge r(h)$,
$\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}$ $=\alpha_{k^{-1}*h}(a1_{h^{-1}*k})\alpha_{k^{-1}}(b1_{k})=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)})\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)})1_{(k^{-1}|f)(f|h)}1_{k^{-1}}\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)})1_{(k^{-1}|f)(f|h)}1_{r(k^{-1}|f)}1_{k^{-1}}\alpha_{k^{-1}}(b1_{k}).$
By Lemma 4.4(ii), we have that $A_{(k^{-1}|f)}=A_{k^{-1}}\cap A_{r(k^{-1}|f)}$, from where it follows that $1_{(k^{-1}|f)}=1_{k^{-1}}1_{r(k^{-1}|f)}$. Therefore,
$\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}$ $=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)})1_{(k^{-1}|f)(f|h)}1_{(k^{-1}|f)}\alpha_{k^{-1}}(b1_{k}).$
Now, by Proposition (i)(ii), we have that
$A_{(k^{-1}|f)(f|h)}\cap A_{(k^{-1}|f)}=\alpha_{(k^{-1}|f)(f|h)}(A_{(h^{-1}|f)(f|k)}\cap A_{(h^{-1}|f)}),$
so $1_{(k^{-1}|f)(f|h)}1_{(k^{-1}|f)}=\alpha_{(k^{-1}|f)(f|h)}(1_{(h^{-1}|f)(f|k)}1_{(h^{-1}|f)})$. Thus,
$\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}$ $=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)})\alpha_{(k^{-1}|f)(f|h)}(1_{(h^{-1}|f)(f|k)}1_{(h^{-1}|f)})\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{(k^{-1}|f)(f|h)}(a1_{(h^{-1}|f)(f|k)}1_{(h^{-1}|f)})\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{(k^{-1}|f)}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})1_{(f|k)})\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{k^{-1}}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})1_{(f|k)})\alpha_{k^{-1}}(b1_{k})$
$=\alpha_{k^{-1}}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})1_{(f|k)}b1_{k})=\alpha_{k^{-1}}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})1_{f}1_{k}b1_{k})$
$=\alpha_{k^{-1}}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})b1_{k})=\psi_{e}(\alpha_{(f|h)}(a1_{(h^{-1}|f)})b)|_{k}.$
Notice that $b\in A_{e}$ implies $\alpha_{(f|h)}(a1_{(h^{-1}|f)})b\in A_{e}$, since $A_{e}$ is an ideal of $A$. Analogously we have that $\psi_{e}(b)\beta_{h}(\psi_{d(h)})\in\psi_{e}(A_{e})$.
The uniqueness now follows by [4, Theorem 2.1]. ∎
Example 4.8.
Resuming Example 3.6 and applying the construction of Theorem 4.7, we obtain $\mathcal{E}_{g}=\mathcal{G},\text{ for all }g\in\mathcal{G}$, from where $F_{g}=\mathcal{F},\text{ for all }g\in\mathcal{G}$.
Thus, $\gamma_{s}(a,b,c,d,f)=(d,c,b,a,f)=\gamma_{s^{-1}}(a,b,c,d,f)$ and $\gamma_{g}=Id_{\mathcal{F}}$, for all $f\in\mathcal{G}_{0}$. Furthermore, notice that $\psi_{r(s)}(ae_{2}+be_{3})=(ae_{2},ae_{2},ae_{2}+be_{3},ae_{2},ae_{2})$ and $\psi_{d(s)}(ae_{2})=\psi_{e}(ae_{2})=(ae_{2},ae_{2},ae_{2},ae_{2},ae_{2})$.
Therefore $B_{s}^{\prime}=\{(ae_{2},ae_{2},ae_{2}+be_{3},ae_{2},ae_{2}):a,b\in R\}$, $B_{s^{-1}}^{\prime}=\{(ae_{2},ae_{2}+be_{3},ae_{2},$ $ae_{2},ae_{2}):a,b\in R\}$ and $B_{e}^{\prime}=Re_{2}(1,1,1,1,1)$. Hence $B^{\prime}=B_{s}^{\prime}+B_{s^{-1}}^{\prime}+B_{e}^{\prime}=\{(ae_{2},ae_{2}+be_{3},ae_{2}+ce_{3},ae_{2},ae_{2}):a,b,c\in R\}$ and then $B^{\prime}\simeq B$ and $B_{g}^{\prime}\simeq B_{g}$, for all $g\in\mathcal{G}$. So $\beta^{\prime}\simeq\beta$.
5. Applications
We conclude by presenting two applications of the results of the previous sections. We first show that the partial skew ordered groupoid ring is Morita equivalent to the skew ordered groupoid ring relative to its globalization. Afterwards, as a consequence of the ESN Theorem, we provide a necessary and sufficient condition for a partial action of an inverse semigroup to have a (unique) minimal globalization.
5.1. A Morita Context
To simplify the notation, in this subsection we will assume that $\varphi_{e}=\iota_{e}:A_{e}\to B_{e}$ is the usual inclusion, so that $A_{e}$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$.
Proposition 5.1.
Let $\alpha=(A_{g},\delta_{g})_{g\in\mathcal{G}}$ be a unital P.O. action of the ordered groupoid $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. Let $R=A\ltimes_{\alpha}^{o}\mathcal{G}$ and $T=B\ltimes_{\beta}^{o}\mathcal{G}$. If $\mathcal{G}_{0}$ is finite,
  1. (i)
    $T1_{R}=\sum_{g\in\mathcal{G}}\overline{\beta_{g}(A_{d(g)})\delta_{g}}$;
  2. (ii)
    $1_{R}T=\sum_{g\in\mathcal{G}}\overline{A_{r(g)}\delta_{g}}$;
  3. (iii)
    $1_{R}T1_{R}=R$;
  4. (iv)
    $T1_{R}T=T$.
Proof.
First notice that the crossed products $A\ltimes_{\alpha}\mathcal{G}$ and $B\ltimes_{\beta}\mathcal{G}$ are associative, so $R$ and $T$ are defined. The proof of (i)-(iii) are similar to the nonordered case [4, Proposition 3.1]. In order to prove (iv), it suffices to show that $T\subseteq 1_{R}T1_{R}$. Since for all $h\in\mathcal{G}$ we have that $B_{h}=\sum_{r(g)\leq r(h)}\beta_{g}(A_{d(g)})$, the result follows from
$\overline{\beta_{g}(a)\delta_{h}}=\overline{\beta_{g}(a)\delta_{(r(g)|h)}}=\overline{\beta_{g}(a)\delta_{g}}\cdot\overline{1_{r(g^{-1}(r(g)|h))}\delta_{g^{-1}(r(g)|h)}}\in(T1_{R})(1_{R}T)=T1_{R}T,$
for all $a\in A_{d(g)}$. ∎
A Morita context is a sextuple $(R,R^{\prime},M,M^{\prime},\varphi,\varphi^{\prime})$, where $R,R^{\prime}$ are unital rings, $M$ is a $(R,R^{\prime})$-bimodule, $M^{\prime}$ is a $(R^{\prime},R)$-bimodule and $\varphi:M\otimes_{R^{\prime}}M^{\prime}\to R$ and $\varphi^{\prime}:M^{\prime}\otimes_{R}M\to R^{\prime}$ are bimodule morphisms such that:
  1. (i)
    $\varphi(x\otimes x^{\prime})y=x\varphi^{\prime}(x^{\prime}\otimes y)$, for all $x,y\in M$ and $x^{\prime}\in M^{\prime}$,
  2. (ii)
    $\varphi^{\prime}(x^{\prime}\otimes x)y^{\prime}=x^{\prime}\varphi(x\otimes y^{\prime})$, for all $x^{\prime},y^{\prime}\in M^{\prime}$ and $x\in M$.
According to [29, Theorems 4.1.4 and 4.1.17], if $\varphi$ and $\varphi^{\prime}$ are surjective, then the categories of left $R$-modules and of left $R^{\prime}$-modules are equivalent and the rings $R$ and $R^{\prime}$ are said to be Morita equivalent. In this case, we also say that the Morita context $(R,R^{\prime},M,M^{\prime},\varphi,\varphi^{\prime})$ is strict.
Theorem 5.2.
The rings $R$ and $T$ as above are Morita equivalent.
Proof.
Consider $M=1_{R}T$ and $N=T1_{R}$. Clearly $M$ is a $(R,T)$-bimodule and $N$ is a $(T,R)$-bimodule. Defining the maps $\varphi:M\otimes_{T}N\to R$ and $\varphi^{\prime}:N\otimes_{R}M\to T$ by $\varphi(m\otimes n)=mn$ and $\varphi^{\prime}(n\otimes m)=nm$, we have that $(R,T,M,N,\varphi,\varphi^{\prime})$ is a strict Morita context by Proposition 5.1. ∎
5.2. Partial Actions of Inverse Semigroups
Let $S$ be an inverse semigroup. According to [7], a partial action of an inverse semigroup $S$ on a ring $A$ is a collection $\alpha=(A_{s},\alpha_{s})_{s\in S}$ of ideals $A_{ss^{-1}}$ of $A$, ideals $A_{s}$ of $A_{ss^{-1}}$ and ring isomorphisms $\alpha_{s}:A_{s^{-1}}\to A_{s}$ such that
  1. (P1)
    $A=\sum_{e\in E(S)}A_{e}$.
  2. (P2)
    $\alpha_{s}(A_{s^{-1}}\cap A_{t})=A_{s}\cap A_{st}$, for all $s,t\in S$.
  3. (P3)
    $\alpha_{s}(\alpha_{t}(a))=\alpha_{st}(a)$, for all $a\in A_{t^{-1}}\cap A_{(st)^{-1}}$, $s,t\in S$.
By [17, Theorem 6.1, Lemma 6.2], we have that the partial actions of an inverse semigroup $S$ on a ring $A$ are in one-to-one correspondence with the strong P.O. actions of an associated inductive groupoid $\mathbb{G}(S)$ on $A$. The ESN Theorem [20, Theorem 4.1.8] also ensures that the global actions of $S$ are in one-to-one correspondence with the global actions of $\mathbb{G}(S)$.
These remarks enable us to translate the theory that we have developed so far to the case of inverse semigroups. More precisely, we have the following result.
Theorem 5.3.
Let $\alpha$ be a partial inverse semigroup action of $S$ on a ring $A$ such that $A_{e}$ is unital for every $e\in E(S)$. Then $\alpha$ has a minimal globalization $\beta$ if and only if $\alpha$ is unital. Moreover, $\beta$ is unique up to isomorphism.
Consider the definition of partial skew inverse semigroup ring given in [7], that is, $A\ltimes_{\alpha}S=\mathcal{L}/\mathcal{N}$, where $\alpha$ is a unital partial action of $S$ on a ring $A$, $\mathcal{L}=\bigoplus_{s\in S}A_{s}\delta_{s}$ with component-wise addition and product given by $(a_{s}\delta_{s})(b_{t}\delta_{t})=\alpha_{s}(\alpha_{s^{-1}}(a_{s})b_{t})\delta_{st}$ and $\mathcal{N}=\langle a_{s}\delta_{s}-a_{s}\delta_{t}:s\leq t\text{ and }a_{s}\in A_{s}\rangle$ is an ideal of $\mathcal{L}$. Then the same argument used in Theorem 5.2 also proves the following result.
Theorem 5.4.
Let $\alpha$ be a unital partial inverse semigroup action of $S$ on a ring $A$. Assume that $E(S)$ is finite. Let $\beta$ be the minimal globalization of $\alpha$, global action of $S$ on a ring $B$. Then $A\ltimes_{\alpha}S$ and $B\ltimes_{\beta}S$ are Morita equivalent.